OpenStudy (anonymous):

When I have something like: Al(OH)3 would I set it up like this: 26.982(15.999+1.008)*3?

OpenStudy (hwyl):

what are you trying to do?

OpenStudy (anonymous):

mole to grams conversions the question is: How many moles are in 98.3 grams of aluminum hydroxide, Al(OH)3

OpenStudy (hwyl):

ah

OpenStudy (hwyl):

how many moles of \(\sf Al(OH)_3 \) do you have?

OpenStudy (anonymous):

thats the question; it only gave grams for me to find that.

OpenStudy (hwyl):

you only have one mole

OpenStudy (hwyl):

if you do not see a number before the name of the compound or molecule then it means there is only one mole.

OpenStudy (hwyl):

\(\sf \huge x~ mole_{Al(OH)_3} \times \frac{grams_{Al(OH)_3}}{1 ~ mole_{Al(OH)_3}} = \)

OpenStudy (hwyl):

do you see how the mole unit for a particular thing is canceled out algebraically?

OpenStudy (hwyl):

\(\large a \times \huge \frac{b}{a} = b\)

OpenStudy (anonymous):

yeah I see where it would be cancelled out. I just need to know how to find the molecular weight of the aluminum hydroxide.

OpenStudy (hwyl):

whatever unit you're trying to convert out of, you need to divide by itself

OpenStudy (hwyl):

molecular weight follows the conservation of mass so if you see Al it means there is 1 Al in there if you see \(\sf (OH)_3\) it means that there are 3 O's and 3 H's; the subscripts tell you how many there are for whatever each element that is enclosed in parenthesis

OpenStudy (hwyl):

usually when there's 1, we do not write it out. it is just like saying a mole of something - it is implicit that the word "a" denotes 1

OpenStudy (hwyl):

you're going to add all the weights of each elements in that molecule/compound

OpenStudy (hwyl):

\(\sf \large Al(OH)_3 = [1 \times (weight_{Al})] + [3 \times (weight_{O}+weight_{H})] \)

OpenStudy (hwyl):

get it?

OpenStudy (anonymous):

so the answer is 78.003 correct?

OpenStudy (hwyl):

I don't know the individual atomic mass units of each elements, that is something you need to reference to a periodic table.

OpenStudy (anonymous):

its in the top

OpenStudy (hwyl):

then start adding and multiplying

OpenStudy (anonymous):

okay well thank you so much for you're help!!

OpenStudy (anonymous):

@hwyl what about when it says: Ca3(PO3)2

OpenStudy (hwyl):

Ca: \(3 \times \sf Ca \) P: \(2 \times \sf P \) O: \(2 \times 3 \sf ~ O \)